OpenStudy (anonymous):

5x + 2y + z = –2 3x + 4y + 3z = 2 –4x – 3y – 3z = 1

OpenStudy (anonymous):

can you use matrices on a TI-83 or in excel?

OpenStudy (anonymous):

If not, here it is by hand: combine two eq's that most easily eliminate a variable, in this case Eq's 2 and 3: 3x + 4y + 3z = 2 –4x – 3y – 3z = 1 Add (vertically) the corresponding terms: -x +y +0 = 3 Then combine one of these with Eq #1 to also eliminate z: multiply all the terms in Eq #1 by 3: 3(5x + 2y + z = –2) --> 15x + 6y +3z = -6 and combine with Eq 3 to eliminate z

OpenStudy (anonymous):

I prefer Excel because there are fewer opportunities for errors